OpenStudy (anonymous):

If integral from 0,2 (2x^3-kx^2+2x)dx=12, then k must be? I've tried working it out several times but I still don' get the right answer.

OpenStudy (anonymous):

\[\int\limits_{}^{?}\]

OpenStudy (anonymous):

\[ int_{0}{2} \]

OpenStudy (anonymous):

\[\int\limits_{0}^{2} (2x^3-kx^2+2k) dx =12\]

OpenStudy (anonymous):

\[\int\limits_{0}^{2} 2x^3 - kx^2 + 2k = 12 \]

OpenStudy (anonymous):

= 2x^4/4 - kx^3/3 +2kx from 2 to 0 , = 12

OpenStudy (anonymous):

2(2^4)/4 - k(2^3)/3 + 2(k)(2) - [ 0 + 0 + 0 ] = 12

OpenStudy (anonymous):

did you get that so far?

OpenStudy (anonymous):

yeah... instead of adding one to the power I was just finding the derivative..

OpenStudy (anonymous):

well we are integrating, correct

OpenStudy (anonymous):

give me a medal

OpenStudy (anonymous):

yeah :) thanks... let me see if i get it now

OpenStudy (anonymous):

Got it :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

hey can u help me in another one real quick? sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\int\limits_{}^{} (\sec^2x)(\tan ^2x)dx=\]

OpenStudy (anonymous):

u = tan x

OpenStudy (anonymous):

du = sec^2 x

OpenStudy (anonymous):

so you do u sub?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so i got integral u^2 du